Solve for p: p(3p+4)=4
Answer by mananth(16949)

p(3p+4)=4
3p^2+4p=4
3p^2+4p-4=0
3p^2+6p-2p-4=0
3p(p+2)-2(p+2)=0
(p+2)(3p-2)=0
p=-2 OR p=2/3
...
m.ananth@hotmail.ca